If f(x)=a^x and f(3)=27, find each value of f(x):
f(-1)
acrown1@umbc.edu

Answer by stanbon(75887)

If f(x)=a^x and f(3)=27, find each value of f(x):
If f(3)=27 then f(3)=a^3=27
Then a=3
So the equation is f(x)=3^x
Then f(-1)=3^(-1)= 1/3
Cheers,
Stan H.